A box with an open top is to be made by taking a rectangular piece of tin 8 by 10 inches and cutting a square of the same size out of each corner and folding up the sides. If the area of the base is to be 24 square inches, what should be the length of the sides of the square be?
Answer by ikleyn(52843)

.
A box with an open top is to be made by taking a rectangular piece of tin 8 by 10 inches
and cutting a square of the same size out of each corner and folding up the sides.
If the area of the base is to be 24 square inches, what should be the length of the sides
of the square be?
~~~~~~~~~~~~~~~~~~~~~~~~~~~

It can be solved in several different ways.


        For example, a  mental solution  is as follows.


The original rectangular piece of tin has dimensions 8 by 10 inches.

The difference between the two dimensions is 10-8 = 2 inches.

After cutting a square of the same size out of each corner and folding up the sides,
the difference between the dimensions of the base will be the same 2 inches.

So, you need to find the dimensions of the rectangular base, given that the area is 
24 square inches, while the difference of the dimensions is 2 inches.


At this point a person, familiar with the multiplication table, will guess the answer 
in 5-6 seconds: the dimensions of the base are 6 by 4 inches.


So, the side of the squares to cut are  (10-6)/2 = 4/2 = 2 inches.


And the problem is just solved mentally to the end this way.



        Algebraic solution  uses equations.


It starts by noticing that if x is the size of the square to cut from each corner,
then, after folding up the sides, the dimensions of the base are (10-2x) inches by (8-2x) inches.


It leads to the equation for the base area

    (10-2x)*(8-2x) = 24  square inches.


So, we should find x from this equation.


Simplify it step by step, but as the first step, divide by sides by 2*2 = 4 to get

    (5-x)*(4-x) = 6,

    20 - 4x - 5x + x^2 = 6

    x^2 - 9x + 14 = 0

    (x-7)*(x-2) = 0     <<<---===  factoring.


Thus two possible roots to the equation are  x= 2  and x= 7.


Bit 7 inches square side is too much to cut (obviously).


It leaves only one remaining possibility x= 2 inches, which is the correct ANSWER.

At this point, the solution is complete,
and from my post, you know two different methods of solving.

Two methods are presented to provide peace in your mind.